# TimeMIL: Advancing Multivariate Time Series Classification via a Time-aware Multiple Instance Learning

### v1

@@ 1 Introduction
Contributions: (i) To the best of our knowledge, we are the first to formally formulate a generic MIL framework for multivariate and multi-class time series classification from an information-theoretic perspective. (ii) We propose a Time-aware MIL pooling based on a tokenized transformer and a novel learnable wavelet positional encoding to model complex patterns within time series. The proposed method outperforms 26 recent state-of-the-art methods in 28 datasets and offers inherent interpretability.
@@ 2 Related Works
Multivariate Time Series Classification. The recent DL methods specifically designed for MTSC can roughly be divided into two categories: (i) CNN/LSTM Hybrid architecture (Karim et al., 2019; Zhang et al., 2020), where LSTM is often used to capture sequential dependencies and CNN is used to capture the inter-variate correlations. (ii) Purely CNN architecture (Ismail Fawaz et al., 2020; Li et al., 2021b; Tang et al., 2022), where long-term dependencies, short-term dependencies, and cross-channel dependencies are claimed to be captured by multiple kernels with varying kernel sizes.
Recently, the authors of (Wu et al., 2023; donghao & wang xue, 2024) offered a model called General Time Series Analysis Framework, to be used in multiple mainstream tasks, including classification, imputation, short-term forecasting, long-term forecasting, and anomaly detection, with simple modifications. Transformer-based models (Zhou et al., 2022; Zhang & Yan, 2023; Nie et al., 2023) and MLP-based models (Zeng et al., 2023; Zhang et al., 2022b; Li et al., 2023c, b) have also been developed and improved over the last few years for this purpose due to their excellent scaling behaviors. Nonetheless, they did not yet fully replace CNN-based models, which continue to exhibit impressive performance (Liu et al., 2022; Wang et al., 2023; donghao & wang xue, 2024).
@@ 3 Method
In this section, we introduce three key components for applying the proposed TimeMIL to MTSC. First, we formulate the MTSC as a MIL problem in section 3.1 and 3.2. Second, we introduce a time-aware MIL pooling to capture the temporal ordering in time series through a wavelet-augmented transformer (Section 3.3). The entire framework of the proposed TimeMIL is depicted in Fig. 2. Third, we introduce interpretable pooling to assess the importance of instance, as discussed in section 3.4. A summary of the proposed framework is presented in Algorithm 1.
@@ 3.1 Problem Formulation
Multivariate Time Series data is typically presented as $\{\boldsymbol{X}_{1},\cdots,\boldsymbol{X}_{n}\}$, where $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$ is a time series with $T$ time points, with each time point $\boldsymbol{x}_{i}^{t}\in\mathbb{R}^{d}$ being a $d$-dimensional vector. It is noteworthy that the time points are shift-variant and ordered. Similar to the standard classification task, the goal of MTSC is to learn a direct mapping from feature space $\mathcal{X}$ to label space $\mathcal{Y}$ using the training data set $\{(\boldsymbol{X}_{1},y_{1}),\cdots,(\boldsymbol{X}_{n},y_{n})\}$, where $y_{i}$ is the label for a single time series.
@@ 3.2 MTSC as a MIL problem
Remark 1 arises from the symmetric (permutation-invariant) property of the MIL pooling function $\sigma$. The function $\sigma$ remains the same for every permutation of the instances within a bag, thereby neglecting the temporal ordering between time points (instances) in time series modeling. This hinders the direct translation of classic MIL into MTSC tasks. To address this limitation, we propose a time-aware MIL pooling in Section 3.3.
@@ 3.3 Time-Aware MIL Pooling for MTSC
From an entropy perspective, understanding permutation-variant properties in time series can be quite insightful. As discussed in Section 3.2, we assume that each time point (instance) $\boldsymbol{x}_{i}^{j}$ in a time series (bag) is a realization of a random variable $\Theta^{j}$ conditioned by a time index $t^{j}$. The resulting bag can be represented as a random variable $\boldsymbol{X}=\{(\Theta^{1}|t^{1}),\cdots,(\Theta^{T}|t^{T})\}$, where $t^{j}\neq t^{k},\forall j\neq k$. Likewise, $\boldsymbol{Y}$ denotes a random variable for bag label. Here, we use notation $\Theta^{j}$ for an instance to distinguish a random variable (often as uppercase) and its realization (often as lowercase). It is fair to construct a general assumption: $p(\Theta^{j}|t^{j})\neq p(\Theta^{j}|t^{k})$, which indicates an instance varies when it is presented in different locations.
@@
where $n_{\operatorname{W}}$ denotes the number of wavelet basis, which is empirically set to 3 in this paper. $\{\boldsymbol{\Psi}_{1},\cdots,\boldsymbol{\Psi}_{n_{\operatorname{W}}}\}$ are learnable wavelet kernels with $\boldsymbol{\Psi}_{j}=\{\psi_{a_{j1},b_{j1}}(t),\cdots,\psi_{a_{jL},b_{jL}}(t)\}$ where $\psi$ is mother wavelet, which is chosen to be the Mexican hat in our experiments. The Gabor-Heisenberg limit (Gabor, 1946), which is the uncertainty principle in the time-frequency version, states that it is impossible to precisely determine both the time and frequency of a signal, simultaneously. In the context of the wavelet transform, this principle implies a trade-off: the higher the resolution needed in time, the lower the resolution becomes in frequency. This implies that the careful selection of time-frequency resolution is crucial for effectively characterizing different signals. Hence, we learn the scaling and translation parameters $\{(a_{j1},b_{j1}),\cdots,(a_{jL},b_{jL})\}$ to form the wavelet basis from mother wavelet. $\Phi(\cdot)$ is the channel-wise wavelet transform, which can be formulated as convolving the input signal with the wavelet kernels:
@@ 3.4 Interpretability
The proposed time-aware MIL is naturally interpretable due to its ability to localize time points of interest within a time series. This is achieved by hacking into the attention map of the transformer layers in the proposed time-aware MIL pooling. The attention map (refer to Eq. 3.3) measures the importance of each instance $\boldsymbol{x}_{i}^{t}$ in the series $\boldsymbol{X}_{i}$ in the MIL pooling:
@@
Algorithm 1 Time-aware MIL (Forward Propagation)
0:  Input sequence $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$. A neural network contains: $f$: feature extractor, $g$: bag-level classifier, $WPE$ bases, and $\boldsymbol{x}^{\operatorname{cls}}$: class token. Two Transfomer blocks.
0:  Predicted label $\hat{y}_{i}$.
4:  for j=1:2 do
7:  end for#Bag-level classification from here
@@ 4.2 Main Experimental Results
The proposed method demonstrates superiority over other recently proposed competing methods in both Group 1 and Group 2 experiments (refer to Table 1 and 2).
@@
Group 2 results. The proposed method also achieved superior performance on the Group 2 datasets compared to other methods. Specifically, our method achieved a 77.7% average accuracy, surpassing the other methods in 5 out of 10 datasets and achieving an average rank of 2.85 out of a total of 17 methods. Notably, the proposed method outperforms the recent state-of-the-art ModernTCN (donghao & wang xue, 2024) by 3.5% in average accuracy and 1.15 in average rank.
@@ 4.4 The effectiveness of weakly supervised learning
Inherent interpretability. The instance-level decision-making mechanism makes the proposed TimeMIL inherently interpretable. Leveraging this, we can obtain the importance score (Eq. 18). TimeMIL accurately localized patterns of interest (i.e., positive instances) in the UEA dataset (Fig. 5; $2^{nd}$ and $3^{rd}$ row) and synthetic dataset (Fig. 5; $1^{st}$ row). This supports our initial hypothesis that time points of interest in positive time series are typically sparse and localized, making the weakly supervised TimeMIL a natural choice for MTSC.
@@ 5 Conclusion
6 Limitation
Currently, our work primarily focuses on modeling the sparsity and locality of time points, and we have not yet extended the interpretability pooling to the channel level. Our work may also be limited by the case that a time point fuses multiple variates that represent potential delayed events or measurements. We envision that a major difference in expanding to the channel level would involve designing a channel-based positional encoding. This will be a topic of exploration in our future work.
7 Impact Statements
This paper presents a novel TimeMIL framework to model the temporal correlation and ordering, leveraging a tokenized transformer with a specialized learnable wavelet positional token. The potential societal consequences of TimeMIL can be summarised three-fold: (i) Theoretical View: Commencing with a viewpoint in weakly supervised learning, we are the first to formally introduce MIL into time series-related tasks. We rigorously examine feasibility from an information-theoretic perspective. In response to the various limitations of MIL in MTSC, we propose corresponding modules to preserve the intrinsic temporal correlation and ordering properties within time series. We draw connections between the MIL and MTSC. This provides an improved justification, relating it to existing MTSC methods. (ii) Flexibility: Our framework significantly expands the MTSC applications. For example, future work involving financial forecasting, predictive maintenance, anomaly detection, healthcare monitoring, financial forecasting, environmental monitoring, and speech and signal processing, among others, can all be incorporated within this framework. (iii) Interpretability: Compared with the previously existing methods, TimeMIL could provide the interpretable interest of the pattern of the network. By visualizing the points of time series that strongly influence predictions, TimeMIL can assist in ensuring model robustness. This can help in identifying vulnerabilities and reducing the risk of adversarial attacks. Meanwhile, Interpretability contributes to the explainability of MTSC, making it easier for researchers, practitioners, and stakeholders to comprehend/validate how and why a model makes specific predictions.
@@ F.2 More detail about the Architecture
As discussed in Section 3.3, we use Nyström Self-attention (Xiong et al., 2021) for accelerating the computation. Specifically, we set the embedding dimension $d_{model}=512$, the number of MHSA heads to $8$. For the Nystrom-based matrix approximation, we set the number of landmark points to 256 and the number of moore-penrose iterations for approximating pseudo-inverse to 6, which is recommended by its original paper. The final classifier consists of two fully connected layers: $\mathbb{R}^{L}\rightarrow\mathbb{R}^{L}\rightarrow\mathbb{R}^{C}$, where $C$ denotes the number of classes. We only feed the class token to the classifier. To facilitate the assumption of TimeMIL that treats MTSC as several one-vs-rest (OvR) binary classifications in the context of MIL, we use the binary cross entropy with one-hot encoding for the sequence label. We also adopt window-based random masking augmentation and a warm-up technique, as discussed below.
@@ Appendix G Additional Experiments
The additional results on all 30 UEA datasets are presented in Table 5. It is observed that classification-specific methods (TapNet, MOS-CNN, TodyNet, and Ours) often perform better. We also include Mamba (Gu & Dao, 2023), the recent state-of-the-art Selective State Spaces model. We implement the vanilla version of mamba with different numbers of blocks {1,2,4,8,10,22}, and the results are presented in Table 6.

### v2

@@ 1 Introduction
Contributions: (i) To the best of our knowledge, we are the first to formally formulate a generic MIL framework for multivariate and multi-class time series classification from an information-theoretic perspective. (ii) We propose a Time-aware MIL pooling based on a tokenized transformer and a novel learnable wavelet positional encoding (WPE) to model complex patterns within time series. The proposed method outperforms 26 recent state-of-the-art methods in 28 datasets and offers inherent interpretability.
@@ 2 Related Works
Multivariate Time Series Classification. The recent DL methods specifically designed for MTSC can roughly be divided into two categories: (i) CNN/LSTM Hybrid architecture (Karim et al., 2019; Zhang et al., 2020), where LSTM is often used to capture sequential dependencies and CNN is used to capture the local features. (ii) Purely CNN architecture (Ismail Fawaz et al., 2020; Li et al., 2021b; Tang et al., 2022), where long-term dependencies, short-term dependencies, and cross-channel dependencies are claimed to be captured by multiple kernels with varying kernel sizes.
Recently, General Time Series Analysis Framework (Wu et al., 2023; donghao & wang xue, 2024) has been also proposed for multiple mainstream tasks, including classification, imputation, short-term forecasting, long-term forecasting, and anomaly detection, with simple modifications. Transformer-based models (Zhou et al., 2022; Zhang & Yan, 2023; Nie et al., 2023) and MLP-based models (Zeng et al., 2023; Zhang et al., 2022b; Li et al., 2023c, b) have also been developed and improved over the last few years for this purpose due to their excellent scaling behaviors. Nonetheless, they did not yet fully replace CNN-based models, which continue to exhibit impressive performance (Liu et al., 2022; Wang et al., 2023; donghao & wang xue, 2024).
@@ 3 Method
In this section, we introduce three key components for applying the proposed TimeMIL to MTSC. First, we formulate the MTSC as a MIL problem in Sec. 3.1 and 3.2. Second, we introduce a time-aware MIL pooling to capture the temporal ordering in time series through a wavelet-augmented transformer (Sec. 3.3). Third, we introduce how to quantify the importance of instances in Sec. 3.4. The entire framework of the proposed TimeMIL is depicted in Fig. 2. A summary of the proposed framework is presented in Algorithm 1.
@@ 3.1 Problem Formulation
Multivariate time series data is typically presented as $\{\boldsymbol{X}_{1},\cdots,\boldsymbol{X}_{n}\}$, where $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$ is a time series contains $T$ time points, with each time point $\boldsymbol{x}_{i}^{t}\in\mathbb{R}^{d}$ being a $d$-dimensional vector. It is noteworthy that the time points are shift-variant and ordered. The goal of MTSC is to learn a direct mapping from feature space $\mathcal{X}$ to label space $\mathcal{Y}$ using the training data $\{(\boldsymbol{X}_{1},y_{1}),\cdots,(\boldsymbol{X}_{n},y_{n})\}$, where $y_{i}$ is the label for each time series.
@@ 3.2 MTSC as A MIL Problem
Remark 1 arises from the symmetric (permutation-invariant) property of the MIL pooling function $\sigma$. The function $\sigma$ remains the same for every permutation of the instances within a bag, thereby neglecting the temporal ordering between time points (instances) in time series modeling. This hinders the direct translation of classic MIL into MTSC tasks. To address this limitation, we propose a time-aware MIL pooling in Sec. 3.3.
Algorithm 1 Time-aware MIL (Forward Propagation)
0:  Input sequence $\boldsymbol{X}_{i}=\left\{\boldsymbol{x}_{i}^{1},\cdots,\boldsymbol{x}_{i}^{T}\right\}$. A neural network contains: $f$: feature extractor, $g$: bag-level classifier, $\operatorname{WPE}$ bases, and class token $\boldsymbol{x}^{\operatorname{cls}}$.
0:  Predicted label $\hat{y}_{i}$.
4:  for j=1:2 do
7:  end for#Bag-level classification from here
@@ 3.3 Time-Aware MIL Pooling for MTSC
From an entropy perspective, understanding permutation-variant properties in time series can be quite insightful. As discussed in Sec. 3.2, we assume that each time point (instance) $\boldsymbol{x}_{i}^{j}$ in a time series (bag) is a realization of a random variable $\Theta^{j}$ conditioned by a time index $t^{j}$. The resulting bag can be represented as a random variable $\boldsymbol{X}=\{(\Theta^{1}|t^{1}),\cdots,(\Theta^{T}|t^{T})\}$, where $t^{j}\neq t^{k},\forall j\neq k$. Likewise, $\boldsymbol{Y}$ denotes a random variable for bag label. Here, we use notation $\Theta^{j}$ for an instance to distinguish a random variable (often as uppercase) and its realization (often as lowercase). It is fair to construct a general assumption: $p(\Theta^{j}|t^{j})\neq p(\Theta^{j}|t^{k})$, which indicates an instance varies when it is presented in different locations.
@@
where $n_{\operatorname{W}}$ denotes the number of wavelet basis, which is empirically set to 3 in this paper. $\{\boldsymbol{\Psi}_{1},\cdots,\boldsymbol{\Psi}_{n_{\operatorname{W}}}\}$ are learnable wavelet kernels with $\boldsymbol{\Psi}_{j}=\{\psi_{a_{j1},b_{j1}}(t),\cdots,\psi_{a_{jL},b_{jL}}(t)\}$ where $\psi$ is mother wavelet, which is chosen to be the Mexican hat in our experiments.
The Gabor-Heisenberg limit (Gabor, 1946), which is the uncertainty principle in the time-frequency version, states that it is impossible to precisely determine both the time and frequency of a signal, simultaneously. In the context of the wavelet transform, this principle implies a trade-off: the higher the resolution needed in time, the lower the resolution becomes in frequency. This implies that the careful selection of time-frequency resolution is crucial for effectively characterizing different signals. Hence, we learn the scaling and translation parameters $\{(a_{j1},b_{j1}),\cdots,(a_{jL},b_{jL})\}$ to form the wavelet basis from mother wavelet. $\Phi(\cdot)$ is the channel-wise wavelet transform, which can be formulated as convolving the input signal with the wavelet kernels:
@@ 3.4 Interpretability
The proposed time-aware MIL is naturally interpretable due to its ability to localize time points of interest within a time series. One way to achieve this is to hack into the attention map of the transformer layers in the proposed time-aware MIL pooling. The attention map (refer to Eq. 3.3) measures the importance of each instance $\boldsymbol{x}_{i}^{t}$ in the series $\boldsymbol{X}_{i}$ in the MIL pooling:
@@ 4.2 Main Experimental Results
The proposed method demonstrates superiority over other recently proposed competing methods in both Group 1 and Group 2 experiments (refer to Table 1 and 2). Please refer to Appendix G for additional results.
@@
Group 2 results. The proposed method also achieved superior performance on the Group 2 datasets compared to other methods. Specifically, our method achieved a 77.7% average accuracy, surpassing the other methods in 5 out of 10 datasets and achieving an average rank of 2.85 out of a total of 17 methods. Remarkably, the proposed method outperforms the recent state-of-the-art ModernTCN (donghao & wang xue, 2024) by 3.5% in average accuracy and 1.15 in average rank.
@@ 4.4 The Effectiveness of Weakly Supervised Learning
Inherent interpretability. The instance-level decision-making mechanism makes the proposed TimeMIL inherently interpretable. Leveraging this, we can obtain the importance score. TimeMIL accurately localized patterns of interest (i.e., positive instances) in the UEA dataset (Fig. 5; $2^{nd}$ and $3^{rd}$ row) and synthetic dataset (Fig. 5; $1^{st}$ row). This supports our initial hypothesis that time points of interest in positive time series are typically sparse and localized, making the weakly supervised TimeMIL a natural choice for MTSC.
@@ 5 Conclusion
Impact Statements
This paper presents a novel TimeMIL framework to model the temporal correlation and ordering, leveraging a tokenized transformer with a specialized learnable wavelet positional token. The potential societal consequences of TimeMIL can be summarized three-fold:
(i) Theoretical View. Commencing with a viewpoint in weakly supervised learning, we are the first to formally introduce MIL into time series-related tasks. We rigorously examine feasibility from an information-theoretic perspective. To address a variety of limitations in applying MIL in MTSC, we propose a time-aware MIL pooling to preserve the intrinsic temporal correlation and ordering properties within time series. In summary, we draw theoretical connections between the MIL and MTSC.
(ii) Applicability. Our framework potentially expands the applications of MTSC to financial forecasting, predictive maintenance, anomaly detection, healthcare monitoring, financial forecasting, environmental monitoring, and speech and signal processing.
(iii) Interpretability. Compared with the previously existing methods, TimeMIL could provide the interpretable interest of the pattern of the network. By visualizing the points of time series that strongly influence predictions, TimeMIL can assist in ensuring model robustness. This can help identify vulnerabilities and reduce the risk of adversarial attacks. Meanwhile, interpretability contributes to the explainability of MTSC, making it easier for researchers, practitioners, and stakeholders to comprehend/validate how and why a model makes specific predictions.
Limitation. We have not yet extended the MIL to consider the cross-channel information. We envision a major difference in expanding the current framework to the multi-channel version of TimeMIL, which could involve designing cross-channel temporal attention and positional encoding. This will be a topic of exploration in future work.
Acknowledgement
This material is based upon the work supported by the National Science Foundation under Grant Number CNS-2204721. It is also supported by our collaborative project with MIT Lincoln Lab under Grant Numbers 2015887 and 7000612889.
@@ F.2 More detail about the Architecture
As discussed in Section 3.3, we use Nyström Self-attention (Xiong et al., 2021) for accelerating the computation. Specifically, we set the embedding dimension $d_{model}=512$, the number of MHSA heads to $8$. For the Nystrom-based matrix approximation, we set the number of landmark points to 256 and the number of moore-penrose iterations for approximating pseudo-inverse to 6, which is recommended by its original paper. The final classifier consists of two fully connected layers: $\mathbb{R}^{L}\rightarrow\mathbb{R}^{L}\rightarrow\mathbb{R}^{C}$, where $C$ denotes the number of classes. We only feed the class token to the classifier. To facilitate the assumption of TimeMIL that treats MTSC as several one-vs-rest (OvR) binary classifications in the context of MIL, we use the binary cross entropy with one-hot encoding for the sequence label. We also adopt window-based random masking augmentation and a warm-up technique, as discussed below. The importance score can be conveniently approximated by using Average-Pooling Based Attention (APBA) proposed by (Zhu et al., 2023).
@@ Appendix G Additional Experiments
The additional results on all 30 UEA datasets are presented in Table 5. It is observed that classification-specific methods (TapNet, MOS-CNN, TodyNet, and Ours) often perform better. We also include Mamba (Gu & Dao, 2023), the recent state-of-the-art Selective State Spaces model. We implement the vanilla version of mamba with different numbers of blocks {1,2,4,8,10,14,22}, and the results are presented in Table 6.